1. A football player of mass 90 kg is running at 10 m/s and catches a football of mass
Answer #1

momentm will be conserved >

90*10 + 0.5*40 = 90.5*v

v = 10.166 m/s

energy lost = 1/2( 90*10^2 + 0.5* 40^2 - 90.5^10.166^2) = 223.523 J